OpenStudy (anonymous):

the second angle of a triangle is 4 times as large as the first, the third angle is 80 degree more than the sum of the other two angles. find the measure of all three angles

OpenStudy (anonymous):

a+4a+80+4a+a=180 10a+80=180 10a=100 a=10 b=4a=40 c=80+a+b=130

OpenStudy (anonymous):

Angle 1: x Angle 2:4x Angle 3: 4x+x+80=5x+80 Since the sum of all the angles of a triangle is 180, you can get x+4x+5x+80=180. then by combining like terms, you get 10x=100, and x=10. So angle 1 is 10, angle 2 is 40, and angle 3 is 130.

OpenStudy (anonymous):

way to wash rinse and repeat